OpenStudy (anonymous):

each back will take 7/8 yards of material. you have 21 yards of material. how many backpacks can she make?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

21/(7/8)=8*3=24

OpenStudy (anonymous):

24